OpenStudy (anonymous):

Solve 2x + 5y = 10 for y.

OpenStudy (anonymous):

2x+5y=10 5y=10-2x y=2-(2/5)x

OpenStudy (saifoo.khan):

sorry, misread the ques!

OpenStudy (anonymous):

:/ /but It Doesnt Have That As A Answer

OpenStudy (saifoo.khan):

\[y=\frac{10-2x}{5}\]

OpenStudy (saifoo.khan):

does it have this?

OpenStudy (anonymous):

No:/

OpenStudy (saifoo.khan):

alright, give the options here.

OpenStudy (anonymous):

y=2

OpenStudy (anonymous):

Wait iGot The answer , Thnks For The Help Though (:

OpenStudy (saifoo.khan):

whts the Answer?